# Palantír: Towards Efficient Super Resolution for Ultra-high-definition Live Streaming

### v1

@@ Abstract.
Neural enhancement through super-resolution (SR) deep neural networks (DNNs) opens up new possibilities for ultra-high-definition live streaming over existing encoding and networking infrastructure. Yet, the heavy SR DNN inference overhead leads to severe deployment challenges. To reduce the overhead, existing systems propose to apply DNN-based SR only on selected anchor frames while upscaling non-anchor frames via the lightweight reusing-based SR approach. However, frame-level scheduling is coarse-grained and fails to deliver optimal efficiency. In this work, we propose Palantír, the first neural-enhanced UHD live streaming system with fine-grained patch-level scheduling. In the presented solutions, two novel techniques are incorporated to make good scheduling decisions for inference overhead optimization and reduce the scheduling latency. Firstly, under the guidance of our pioneering and theoretical analysis, Palantír constructs a directed acyclic graph (DAG) for lightweight yet accurate quality estimation under any possible anchor patch set. Secondly, to further optimize the scheduling latency, Palantír improves parallelizability by refactoring the computation subprocedure of the estimation process into a sparse matrix-matrix multiplication operation. The evaluation results suggest that Palantír incurs a negligible scheduling latency accounting for less than 5.7% of the end-to-end latency requirement. When compared to the state-of-the-art real-time frame-level scheduling strategy, Palantír reduces the energy overhead of SR-integrated mobile clients by 38.1% at most (and 22.4% on average) and the monetary costs of cloud-based SR by 80.1% at most (and 38.4% on average). The source code is available at https://palantir-sr.github.io.
@@ 1. Introduction
Ultra-high-definition (UHD) videos such as 4K and 8K videos are expected to form a huge market worth more than \$1 trillion in the following few years (Global Market Insights Inc., 2024). More and more UHD live-streaming applications are deployed to revolutionize many aspects of our society. For example, the UHD live streaming of major sports events such as the Olympics (International Olympic Committee, 2024, 2022) creates unprecedentedly immersive experiences for audiences. Besides, with the real-time UHD video from inspection and surveillance cameras, human operators can have a precise understanding of the spot and remotely take immediate actions to prevent emergent accidents (Coastal Safety Group, 2021; Flyability, 2024).
@@ 4.1. Problems with Existing Methods
    Second, we argue that the types of sub-frame encoding units such as macroblocks (MBs) do not implicitly imply their degrees of reference. Here we conduct a preliminary experiment on the 480p version of the first benchmark video used in the evaluation part (Sec. 6) to support our argument. We separately measure the distribution of degrees of reference among different types of frames and macroblocks (including prioritized frame types, unprioritized normal frames, intra-coded MBs, and inter-coded MBs) The degree of reference for a given frame (or MB) is quantitatively defined as $\frac{num\_references}{resolution}$, where $num\_references$ denotes the number of pixels that refer to the given frame (or MB) for inter coding and $resolution$ denotes the number of pixels in the given frame (or MB). As shown in Fig. 4, the degrees of references for keyframes and AltRefs are significantly greater than those of normal frames, while the distribution range of degrees of references for intra-coded MBs is similar to that for inter-coded MBs. We further use the common language effect size (CLES) (McGraw and Wong, 1992) to quantize the results. The CLES is defined as the probability that a value randomly sampled from one distribution will be greater than that from another distribution. The CLES reaches 97.4% (close to the best case of 100%) between prioritized frame types and normal frames but is as low as 58.9% (close to the worst case of 50%) between different MB types. Therefore, type-based prioritization is helpful in frame-level scheduling but hardly applies to patch-level scheduling.
@@ 6.1. Experimental Setup
Implementation.   We develop our decoder based on the open-source SR decoder in NEMO (Yeo et al., 2020). We incorporate two novel modes into the decoder. The first is to obtain the data required for graph construction (as introduced in Sec. 4.3). The second is to take both an LR video and a corresponding cache profile as input, and then upscale patches by either SR DNNs or reusing-based SR based on the cache profile.
@@
Baselines.   We use two baselines in this part. The first is the NeuroScaler baseline, which uses the algorithm in NeuroScaler to select the anchor frame set. The second is the Key+Uniform baseline, which selects all the patches in the keyframe and equally spaced patches in the remaining frames as anchor patches.
@@ 6.2. Anchor Effectiveness
Quality Gain.   We compute the peak-signal-to-noise-ration (PSNR) between the SR video and the original HR video to quantify the effectiveness of an anchor set. To make a fair comparison, we keep the total size of the anchor regions selected by different methods the same. In other words, we compare the quality gain under the $m$-ary anchor frame set selected by the NeuroScaler baseline with that under the $(15\cdot m)$-ary anchor patch set selected by Palantír or the Key+Uniform baseline. We empirically limit $m$ to not be greater than $3$ since: (1) $m=3$ can deliver quality gains that are comparable to the setting of applying DNN-based SR on all frames; (2) further increasing the value of $m$ leads to a limited benefit yet a significant overhead.
(a) Setting: 1 anchor frame / 15 anchor patches.
(b) Setting: 2 anchor frames / 30 anchor patches.
(c) Setting: 3 anchor frames / 45 anchor patches.
Figure 8. A comparison of anchor effectiveness.
As shown in Fig. 8, our system consistently outperforms the two baselines with its ability to identify beneficial patches. Palantír boosts the quality gain of neural enhancement by 3.7 times at most and 1.4 times on average than the NeuroScaler baseline, or by 3.7 times at most and 1.7 times on average than the Key+Uniform baseline. We also observe that the Key+Uniform baseline even falls behind the NeuroScaler baseline, so the effect of fine-grained scheduling heavily depends on the anchor selection method.

### v2

@@ Abstract.
Neural enhancement through super-resolution (SR) deep neural networks (DNNs) opens up new possibilities for ultra-high-definition (UHD) live streaming over existing encoding and networking infrastructure. Yet, the heavy SR DNN inference overhead leads to severe deployment challenges. To reduce the overhead, existing systems propose to apply DNN-based SR only on carefully selected anchor frames while upscaling non-anchor frames via the lightweight reusing-based SR approach. However, frame-level scheduling is coarse-grained and fails to deliver optimal efficiency. In this work, we propose Palantír, the first neural-enhanced UHD live streaming system with fine-grained patch-level scheduling. Two novel techniques are incorporated into Palantír to select the most beneficial anchor patches and support latency-sensitive UHD live streaming applications. Firstly, under the guidance of our pioneering and theoretical analysis, Palantír constructs a directed acyclic graph (DAG) for lightweight yet accurate SR quality estimation under any possible anchor patch set. Secondly, to further optimize the scheduling latency, Palantír improves parallelizability by refactoring the computation subprocedure of the estimation process into a sparse matrix-matrix multiplication operation.
The evaluation results suggest that Palantír incurs a negligible scheduling latency accounting for less than 5.7% of the end-to-end latency requirement. When compared to the naive method of applying DNN-based SR on all the frames, Palantír can reduce the SR DNN inference overhead by 20 times (or 60 times) while preserving 54.0-82.6% (or 32.8-64.0%) of the quality gain. When compared to the state-of-the-art real-time frame-level scheduling strategy, Palantír can reduce the SR DNN inference overhead by 80.1% at most (and 38.4% on average) without sacrificing the video quality.
@@ 1. Introduction
UHD videos such as 4K and 8K videos are expected to form a huge market worth more than \$1 trillion in the following few years (Global Market Insights Inc., 2024). More and more UHD live-streaming applications are deployed to revolutionize many aspects of our society. For example, the UHD live streaming of major sports events such as the Olympics (International Olympic Committee, 2024, 2022) creates unprecedentedly immersive experiences for audiences. Besides, with the real-time UHD video from inspection and surveillance cameras, human operators can have a precise understanding of the spot and remotely take immediate actions to prevent emergent accidents (Coastal Safety Group, 2021; Flyability, 2024).
@@ 4.1. Problems with Existing Methods
    Second, we argue that the types of sub-frame encoding units such as macroblocks (MBs) do not implicitly imply their degrees of reference. Here we conduct a preliminary experiment on the 480p version of the first benchmark video used in the evaluation part (Sec. 6) to support our argument. We separately measure the distribution of degrees of reference among different types of frames and macroblocks (including prioritized frame types, unprioritized normal frames, intra-coded MBs, and inter-coded MBs). The degree of reference for a given frame (or MB) is quantitatively defined as $\frac{num\_references}{resolution}$, where $num\_references$ denotes the number of pixels that refer to the given frame (or MB) for inter coding and $resolution$ denotes the number of pixels in the given frame (or MB). As shown in Fig. 4, the degrees of references for keyframes and AltRefs are significantly greater than those of normal frames, while the distribution range of degrees of references for intra-coded MBs is similar to that for inter-coded MBs. We further use the common language effect size (CLES) (McGraw and Wong, 1992) to quantize the results. The CLES is defined as the probability that a value randomly sampled from one distribution will be greater than that from another distribution. The CLES reaches 97.4% (close to the best case of 100%) between prioritized frame types and normal frames but is as low as 58.9% (close to the worst case of 50%) between different MB types. Therefore, type-based prioritization is helpful in frame-level scheduling but hardly applies to patch-level scheduling.
@@ 6.1. Experimental Setup
Implementation.   We develop our decoder based on the open-source SR decoder in NEMO (Yeo et al., 2020). We incorporate two novel modes into the decoder. The first is to obtain the data required for graph construction (as introduced in Sec. 4.3). The second is to take both an LR video and a corresponding cache profile as input, and then upscale patches by either SR DNNs or reusing-based SR based on the cache profile. The source code is available at https://palantir-sr.github.io.
@@
Baselines.   We use three baselines in this part. The first is the Per-frame baseline, which applies DNN-based SR on all the frames. The second is the NeuroScaler baseline, which uses the algorithm in NeuroScaler (Yeo et al., 2022) to select the anchor frame set. The third is the Key+Uniform baseline, which selects all the patches in the keyframe and equally spaced patches in the remaining frames as anchor patches.
@@
(a) Setting: $m=1$.
(b) Setting: $m=2$.
(c) Setting: $m=3$.
Figure 8. A comparison of anchor effectiveness.
@@ 6.2. Anchor Effectiveness
Quality Gain.   We compute the peak-signal-to-noise-ration (PSNR) between the SR video and the original HR video to quantify the effectiveness of an anchor set. To make a fair comparison, we keep the total sizes of the anchor regions the same, i.e., compare the quality gain under the $m$-ary anchor frame set selected by the NeuroScaler baseline with that under the $(15\cdot m)$-ary anchor patch set selected by Palantír or the Key+Uniform baseline. The only exception here is the Per-frame baseline, where all the frames are always treated as anchors and the size of the anchor regions is thus always larger than other methods. Furthermore, we empirically limit $m$ to not be greater than $3$ since: (1) $m=3$ can deliver quality gains that are comparable to the setting of applying DNN-based SR on all frames; (2) further increasing the value of $m$ leads to a limited benefit yet a significant overhead.
The results are shown in Fig. 8, from which we have three observations: (1) Palantír consistently outperforms the NeuroScaler baseline and the Key+Uniform baseline with its ability to identify beneficial patches. Palantír boosts the quality gain of neural enhancement by 3.7 times at most and 1.4 times on average than the NeuroScaler baseline, or by 3.7 times at most and 1.7 times on average than the Key+Uniform baseline. (2) The fine-grained scheduling-based Key+Uniform baseline even falls behind the coarse-grained scheduling-based NeuroScaler baseline, so the effect of fine-grained scheduling heavily depends on the anchor selection method. (3) Palantír reduces the SR DNN inference overhead by 20 times with $m=3$ (or 60 times with $m=1$) while compared to the Per-frame baseline. With such a remarkable overhead reduction, Palantír still preserves 54.0-82.6% (or 32.8-64.0%) of the quality gain of the Per-frame baseline.